# Learning to Receive Help: Intervention-Aware Concept Embedding Models

### v1

@@ 4.1 Task and Concept Performance in the Absence of Interventions (Q1)
Table 1: Task predictive performance (accuracy or AUC in %) and mean concept AUC (%) across all tasks and baselines. Values within one standard deviation from the highest baseline are emphasised. CelebA’s task accuracies are relatively low across all baselines given its inherent difficulty (256 highly imbalanced classes). Nevertheless, we highlight these results match previous observations Espinosa Zarlenga et al. (2022).
@@ 4.2 Intervention Performance (Q2)
Given that a core appeal of CBM-like models is their receptiveness to interventions, we now explore how test-time interventions affect IntCEM’s task performance compared to other models. We first study whether IntCEM’s training procedure preconditions it to be more receptive to test-time interventions by evaluating the task accuracy of all baselines after receiving random concept interventions. Then, we investigate whether IntCEM’s competitive advantage holds when exposing our baselines to state-of-the-art intervention policies. Below, we report our findings by showing each model’s task accuracy as we intervene, following Koh et al. (2020), on an increasing number of groups of mutually exclusive concepts (e.g., “white wing” and “black wing”).
@@
We explore IntCEM’s receptiveness to different test-time intervention policies by computing its test performance while intervening following: (1) a Random intervention policy, where the next concept is selected, uniformly at random, from the set of unknown concepts, (2) the Uncertainty of Concept Prediction (UCP) Shin et al. (2023) policy, where the next concept is selected by choosing the concept $c_{i}$ whose predicted probability $\hat{p}_{i}$ has the highest uncertainty (measured by $1/|\hat{p}_{i}-0.5|$), (3) the Cooperative Policy (CooP) Chauhan et al. (2022), where we select the concept that maximises a linear combination of its predicted uncertainty (akin to UCP) and the expected change in the predicted label’s probability when intervening on that concept, (4) Concept Validation Accuracy (CVA) static policy, where concepts are selected using a fixed order starting with those whose validation errors are the highest as done by Koh et al. (2020), (5) the Concept Validation Improvement (CVI) Chauhan et al. (2022) static policy, where concepts are selected using a fixed order starting with those concepts that, on average, lead to the biggest improvements in validation accuracy when intervened on, and finally (6) an oracle Skyline policy, which selects $c_{*}$ on every step, indicating an upper bound for all greedy policies. For details on each policy baseline, including how their hyperparameters are selected, see Appendix 6.7.
We demonstrate in Figure 4 that IntCEM consistently outperforms CEMs regardless of the test-time intervention policy used, with CooP and UCP consistently yielding superior results. Further, we uncover that intervening on IntCEM with a Random policy at test-time outperforms the theoretical best performance attainable by a CEM (Skyline) under many interventions, while its best-performing policy outperforms CEM’s Skyline after very few interventions. These results suggest that IntCEM’s train-time conditioning not only results in IntCEMs being more receptive to test-time interventions but may lead to significant changes in an IntCEM’s theoretical-optimal intervention performance. We include analyses on our remaining datasets, revealing similar trends, in Appendix 6.7.
Figure 4: Task accuracy of IntCEMs and CEMs on CUB and CelebA when intervening with different test-time policies. We show similar improvements of IntCEMs over CBMs in Appendix 6.7.
@@ 4.3 Studying IntCEM’s Intervention Policy (Q3)
Finally, we explore IntCEM’s intervention policy by evaluating test-time interventions predicted by $\psi$. To understand whether there is a benefit of learning $\psi$ end-to-end (i.e., in conjunction with the rest of model training), rather than learning it post-hoc after training an IntCEM, we compare it against a Behavioural Cloning (BC) Bain and Sammut (1995) policy “BC-Skyline” trained on demonstrations of Skyline applied to a trained IntCEM. Furthermore, to explore how $\psi$’s trajectories at train-time help improve IntCEM’s receptiveness to test-time interventions, we study an IntCEM trained by sampling trajectories uniformly at random (“IntCEM no $\psi$”). Our results in Figure 5 suggest that: (1) IntCEM’s learnt policy $\psi$ leads to test-time interventions that are as good or better than CooP’s, yet avoid CooP’s computational cost (up to $\sim$2x faster as seen in Appendix 6.5), (2) learning $\psi$ end-to-end during training yields a better policy than one learnt through BC after the IntCEM is trained, as seen in the BC policy’s lower performance; and (3) using a learnable intervention policy at train-time results in a significant boost of test-time performance. This suggests that part of IntCEM’s receptiveness to interventions, even when using a random test-time policy, can be partially attributed to learning and using $\psi$ during training. We show some qualitative examples of rollouts of $\psi$ in the CUB dataset in Figure 6.
@@ 5 Discussion and Conclusion
Previous work Mahinpei et al. (2021); Havasi et al. (2022); Espinosa Zarlenga et al. (2023) has shown that CBMs are prone to encoding unnecessary information in their learnt concept representations. This phenomenon, called concept leakage, may lead to less interpretable concept representations Mahinpei et al. (2021) and detrimental concept interventions in CBMs Espinosa Zarlenga et al. (2023). Given that interventions in IntCEMs involve swapping a concept’s predicted embeddings rather than overwriting an activation in the bottleneck, such interventions are distinct from those in CBMs as they enable leaked information to be exploited after an intervention is performed. This means our loss function may incentivise IntCEM to exploit this mechanism to improve a model’s receptiveness to interventions. In Appendix 6.10 we explore this hypothesis and find that we can detect more leakage in IntCEM’s concept representations than in those learnt by CBMs and CEMs. This suggests that, contrary to common assumptions, leakage may be a healthy byproduct of more expressive concept representations in models that accommodate such expressivity. Nevertheless, we believe further work is needed to understand the consequences of this leakage.
@@
To the best of our knowledge, we are the first to frame learning to receive concept interventions as a joint optimisation problem where we simultaneously learn concepts, downstream labels, and policies. IntCEMs offer a principled way to prepare concept-based models for interventions at test-time, without sacrificing performance in the absence of interventions. Here, we focused on CEMs as our base model. However, our method can be extended to traditional CBMs (see Appendix 6.8), and future work may consider extending it to more recent concept-based models such as post-hoc CBMs Yuksekgonul et al. (2023), label-free CBMs Oikarinen et al. (2023), and probabilistic CBMs Kim et al. (2023).
Furthermore, we note some limitations. First, IntCEM requires hyperparameter tuning for $\lambda_{\text{roll}}$ and $\lambda_{\text{concept}}$. While our ablations in Appendix 6.6 suggest that an IntCEM’s performance gains extend across multiple values of $\lambda_{\text{roll}}$, users still need to tune such parameters for maximising IntCEM’s utility. These challenges are exacerbated by the computational costs of training IntCEMs due to their train-time trajectory sampling (see Appendix 6.5). However, such overhead gets amortised over time given $\psi$’s efficiency over competing policies. Further, we recognise our training procedure renders IntCEMs more receptive to any form of intervention, including adversarial interactions (see Appendix 6.9). This raises a potential societal concern when deciding how our proposed model is deployed. Future work may explore such issues by incorporating error-correcting mechanisms or by considering intervention-time human uncertainty Collins et al. (2023).

### v3

@@ 4.1 Task and Concept Performance in the Absence of Interventions (Q1)
Table 1: Task predictive performance (accuracy or AUC in %) and mean concept AUC (%) across all tasks and baselines. Values within one standard deviation from the highest baseline are emphasised. CelebA’s task accuracies are relatively low across all baselines given its inherent difficulty (256 highly imbalanced classes). Nevertheless, we highlight these results match previous observations [5].
@@ 4.2 Intervention Performance (Q2)
Given that a core appeal of CBM-like models is their receptiveness to interventions, we now explore how test-time interventions affect IntCEM’s task performance compared to other models. We first study whether IntCEM’s training procedure preconditions it to be more receptive to test-time interventions by evaluating the task accuracy of all baselines after receiving random concept interventions. Then, we investigate whether IntCEM’s competitive advantage holds when exposing our baselines to state-of-the-art intervention policies. Below, we report our findings by showing each model’s task accuracy as we intervene, following [4], on an increasing number of groups of mutually exclusive concepts (e.g., “white wing” and “black wing”).
@@
We explore IntCEM’s receptiveness to different test-time intervention policies by computing its test performance while intervening following: (1) a Random intervention policy, where the next concept is selected, uniformly at random, from the set of unknown concepts, (2) the Uncertainty of Concept Prediction (UCP) [8] policy, where the next concept is selected by choosing the concept $c_{i}$ whose predicted probability $\hat{p}_{i}$ has the highest uncertainty (measured by $1/|\hat{p}_{i}-0.5|$), (3) the Cooperative Policy (CooP) [9], where we select the concept that maximises a linear combination of its predicted uncertainty (akin to UCP) and the expected change in the predicted label’s probability when intervening on that concept, (4) Concept Validation Accuracy (CVA) static policy, where concepts are selected using a fixed order starting with those whose validation errors are the highest as done by Koh et al. [4], (5) the Concept Validation Improvement (CVI) [9] static policy, where concepts are selected using a fixed order starting with those concepts that, on average, lead to the biggest improvements in validation accuracy when intervened on, and finally (6) an oracle Skyline policy, which selects $c_{*}$ on every step, indicating an upper bound for all greedy policies. For details on each policy baseline, including how their hyperparameters are selected, see Appendix A.7.
We demonstrate in Figure 4 that IntCEM consistently outperforms CEMs regardless of the test-time intervention policy used, with CooP and UCP consistently yielding superior results. Further, we uncover that intervening on IntCEM with a Random policy at test-time outperforms the theoretical best performance attainable by a CEM (Skyline) under many interventions, while its best-performing policy outperforms CEM’s Skyline after very few interventions. These results suggest that IntCEM’s train-time conditioning not only results in IntCEMs being more receptive to test-time interventions but may lead to significant changes in an IntCEM’s theoretical-optimal intervention performance. We include analyses on our remaining datasets, revealing similar trends, in Appendix A.7.
Figure 4: Task accuracy of IntCEMs and CEMs on CUB and CelebA when intervening with different test-time policies. We show similar improvements of IntCEMs over CBMs in Appendix A.7.
@@ 4.3 Studying IntCEM’s Intervention Policy (Q3)
Finally, we explore IntCEM’s intervention policy by evaluating test-time interventions predicted by $\psi$. To understand whether there is a benefit of learning $\psi$ end-to-end (i.e., in conjunction with the rest of model training), rather than learning it post-hoc after training an IntCEM, we compare it against a Behavioural Cloning (BC) [26] policy “BC-Skyline” trained on demonstrations of Skyline applied to a trained IntCEM. Furthermore, to explore how $\psi$’s trajectories at train-time help improve IntCEM’s receptiveness to test-time interventions, we study an IntCEM trained by sampling trajectories uniformly at random (“IntCEM no $\psi$”). Our results in Figure 5 suggest that: (1) IntCEM’s learnt policy $\psi$ leads to test-time interventions that are as good or better than CooP’s, yet avoid CooP’s computational cost (up to $\sim$2x faster as seen in Appendix A.5), (2) learning $\psi$ end-to-end during training yields a better policy than one learnt through BC after the IntCEM is trained, as seen in the BC policy’s lower performance; and (3) using a learnable intervention policy at train-time results in a significant boost of test-time performance. This suggests that part of IntCEM’s receptiveness to interventions, even when using a random test-time policy, can be partially attributed to learning and using $\psi$ during training. We show some qualitative examples of rollouts of $\psi$ in the CUB dataset in Figure 6.
@@ 5 Discussion and Conclusion
Previous work [36, 6, 10] has shown that CBMs are prone to encoding unnecessary information in their learnt concept representations. This phenomenon, called concept leakage, may lead to less interpretable concept representations [36] and detrimental concept interventions in CBMs [10]. Given that interventions in IntCEMs involve swapping a concept’s predicted embeddings rather than overwriting an activation in the bottleneck, such interventions are distinct from those in CBMs as they enable leaked information to be exploited after an intervention is performed. This means our loss function may incentivise IntCEM to exploit this mechanism to improve a model’s receptiveness to interventions. In Appendix A.10 we explore this hypothesis and find that we can detect more leakage in IntCEM’s concept representations than in those learnt by CBMs and CEMs. This suggests that, contrary to common assumptions, leakage may be a healthy byproduct of more expressive concept representations in models that accommodate such expressivity. Nevertheless, we believe further work is needed to understand the consequences of this leakage.
@@
To the best of our knowledge, we are the first to frame learning to receive concept interventions as a joint optimisation problem where we simultaneously learn concepts, downstream labels, and policies. IntCEMs offer a principled way to prepare concept-based models for interventions at test-time, without sacrificing performance in the absence of interventions. Here, we focused on CEMs as our base model. However, our method can be extended to traditional CBMs (see Appendix A.8), and future work may consider extending it to more recent concept-based models such as post-hoc CBMs [37], label-free CBMs [38], and probabilistic CBMs [39].
Furthermore, we note some limitations. First, IntCEM requires hyperparameter tuning for $\lambda_{\text{roll}}$ and $\lambda_{\text{concept}}$. While our ablations in Appendix A.6 suggest that an IntCEM’s performance gains extend across multiple values of $\lambda_{\text{roll}}$, users still need to tune such parameters for maximising IntCEM’s utility. These challenges are exacerbated by the computational costs of training IntCEMs due to their train-time trajectory sampling (see Appendix A.5). However, such overhead gets amortised over time given $\psi$’s efficiency over competing policies. Further, we recognise our training procedure renders IntCEMs more receptive to any form of intervention, including adversarial interactions (see Appendix A.9). This raises a potential societal concern when deciding how our proposed model is deployed. Future work may explore such issues by incorporating error-correcting mechanisms or by considering intervention-time human uncertainty [29].